# Uni-3DAR: Unified 3D Generation and Understanding via Autoregression on Compressed Spatial Tokens

### v2

@@ Abstract
Recent advancements in large language models and their multi-modal extensions have demonstrated the effectiveness of unifying generation and understanding through autoregressive next-token prediction. However, despite the critical role of 3D structural generation and understanding (3D GU) in AI for science, these tasks have largely evolved independently, with autoregressive methods remaining underexplored. To bridge this gap, we introduce Uni-3DAR, a unified framework that seamlessly integrates 3D GU tasks via autoregressive prediction. At its core, Uni-3DAR employs a novel hierarchical tokenization that compresses 3D space using an octree, leveraging the inherent sparsity of 3D structures. It then applies an additional tokenization for fine-grained structural details, capturing key attributes such as atom types and precise spatial coordinates in microscopic 3D structures. We further propose two optimizations to enhance efficiency and effectiveness. The first is a two-level subtree compression strategy, which reduces the octree token sequence by up to 8x. The second is a masked next-token prediction mechanism tailored for dynamically varying token positions, significantly boosting model performance. By combining these strategies, Uni-3DAR successfully unifies diverse 3D GU tasks within a single autoregressive framework. Extensive experiments across multiple microscopic 3D GU tasks, including molecules, proteins, polymers, and crystals, validate its effectiveness and versatility. Notably, Uni-3DAR surpasses previous state-of-the-art diffusion models by a substantial margin, achieving up to 256% relative improvement while delivering inference speeds up to 21.8x faster. The code is publicly available at https://github.com/dptech-corp/Uni-3DAR.
@@ 1 Introduction
Figure 1: Overview of the Uni-3DAR framework. It supports diverse 3D structures from microscopic to macroscopic scales via our proposed tokenizer. Its autoregressive modeling enables both 3D structure generation and understanding within a single model.
Recent successes in large language models (LLMs) demonstrate that unifying generation and understanding is essential for generalist models [1, 2]. In deep learning, generation refers to synthesizing new data, while understanding is reflected in tasks like classification and regression. LLMs achieve this unification via autoregressive next-token prediction. Moreover, this paradigm has been extended to multi-modal data (e.g., speech, images, and video) by tokenizing them within a unified next-token prediction framework, ultimately leading to a unified multi-modal large language model [3, 4, 5].
In AI for science, accurately modeling three-dimensional (3D) structures is critical, as even minor spatial differences can profoundly impact reactivity and functionality. Tasks such as protein folding [6], molecular docking [7], and property prediction [8, 9, 10] heavily rely on capturing these spatial interactions. However, existing approaches for 3D generation and understanding have typically been developed independently. Generation predominantly employs diffusion models to synthesize structures from noise [11, 12], while understanding usually relies on unsupervised, BERT-style pretraining [13, 8]. Furthermore, previous models are often tailored for specific types of 3D structures. For example, models designed for crystal structures cannot be directly applied to proteins. Such separate developments limit overall effectiveness. Moreover, although autoregressive models have demonstrated great success in natural language and image domains, their application to 3D structural data remains largely unexplored. This paper aims to bridge this gap by proposing an autoregressive prediction framework that unifies 3D structural generation and understanding across diverse 3D structural data—including molecules, proteins, polymers, and crystals  [^1] Although we primarily use microscopic 3D data in our experiments to illustrate and validate our method, the proposed approach can be seamlessly extended to various types of sparse 3D data..
One critical challenge in achieving this unified framework is designing an effective tokenization strategy that converts a 3D structure into a 1D sequence of tokens. Existing approaches broadly categorized into two types, each with significant limitations. Point-based (or atom-based in microscopic structures) tokenization leverages the sparsity of 3D structures by sequentially predicting points’ coordinates and properties (e.g., atom types) [14, 15]. However, it struggles to simultaneously generate 3D coordinates, properties, and the corresponding prediction order. Grid-based tokenization discretizes 3D space into voxel grids [16, 17, 18], enabling the application of image-based token prediction techniques. However, this approach suffers from severe computational inefficiency due to the cubic growth in token count. For instance, even small molecules may require tens of thousands of tokens [18], which makes such methods impractical for real-world applications. In summary, point-based methods exploit the inherent sparsity of 3D structures but fail to capture full spatial context, while grid-based methods retain complete spatial information at the cost of efficiency. This trade-off underscores the need for a more balanced and scalable approach to 3D structural tokenization.
Figure 2: Overview of the proposed tokenization (illustrated in 2D using a quadtree for clarity). (a) Adaptive coarse-to-fine subdivision of grid cells, where darker nodes indicate non-empty cells that can be further partitioned. (b) This partitioning process constructs an octree, providing a lossless compression of the full-size 3D grid. (c) Uni-3DAR’s tokenization consists of two components: hierarchical spatial compression via an octree and fine-grained structural tokenization. Each node’s position is determined by its tree level and cell center. (d) The proposed 2-level subtree compression further reduces the octree tokens by up to 8x (4x in the illustrated quadtree).
To address these limitations, we propose a hierarchical tokenization method that losslessly and efficiently compresses the full-size 3D grid. As illustrated in Fig. 2 (a) and (b), we first construct an octree hierarchically based on data sparsity and a predefined maximum depth. We then introduce a fine-grained structural tokenization that encodes essential details—such as atomic types and precise coordinates for microscopic 3D structures. Concatenating these tokens level by level produces a hierarchical, coarse-to-fine 1D token sequence that effectively represents the 3D structure (Fig. 2 (c)). Compared to point-based methods, our hierarchical tokenization preserves approximate 3D positional context from preceding levels, thereby improving prediction accuracy. In contrast to full-size grid-based approaches, it efficiently compresses 3D space and avoids cubic token growth, making it significantly more scalable for real-world applications.
Building on this hierarchical tokenization, we employ autoregressive transformer models [19] to unify 3D structural generation and understanding. We call this model Uni-3DAR, with two further optimizations:
1. 2-Level Subtree Compression: Rather than assigning each tree node an individual token, we compress a 2-level subtree (8 subcells) into a single token (Fig. 2 (d)). As each subcell is classified binary (empty or not), grouping 8 subcells results in $2562^{8}=$ 256 distinct states, reducing the sequence length approximately 8x and converting 8 binary classifications into one 256-class task.
@@
2. Masked Next-Token Prediction: Unlike text or image autoregressive models, where the next token’s position is fixed, our hierarchical tokenization produces dynamic token positions due to inherent sparsity. Without an explicitly provided next position, the model cannot reliably predict the next token. Simply appending the next position to the current token did not yield satisfactory results. To address this, we propose a masked next-token prediction strategy that combines masked prediction with next-token prediction. Specifically, as illustrated in Fig. 3 (a), we duplicate each token so that it appears twice with the same token position, and then replace the first copy with a “[MASK]” token. Based on this duplicated sequence, we still use next-token prediction, but exclusively on the masked token. Notably, this setup ensures that prediction is performed using the correct positional information for the intended next token, effectively addressing the challenge of dynamic token positions. Although this approach doubles the sequence length, it leads to significant performance gains.
By employing duplicated masked tokens, Uni-3DAR seamlessly unifies 3D structural generation and understanding tasks within a single framework (Fig. 3 (b)). In this setup, generation leverages masked tokens, token-level (atom-level) understanding operates on fine-grained tokens, and structure-level tasks focus on the “[EoS]” (End of Structure) tokens, ensuring distinct roles without mutual interference. Additionally, the autoregressive architecture inherently supports conditional and multi-frame structural generation.
@@
Extensive experiments across diverse tasks—including molecular and crystal generation, protein pocket prediction, molecular docking, and molecular pretraining—demonstrate Uni-3DAR’s superior or competitive performance compared to existing methods. Notably, Uni-3DAR consistently outperforms diffusion-based generation models, underscoring its strong generative capabilities. Further evaluations confirm the benefits of unifying generation and understanding, highlight the contributions of each component via ablation studies, and reveal significant efficiency advantages in inference speed. Overall, these results validate the effectiveness of the proposed Uni-3DAR framework.
@@ 2.1 Dynamic Coarse-to-Fine Tokenization for 3D Structures
3D structures inherently exhibit sparsity. At the microscopic scale, most space is empty except for a few scattered atoms; similarly, at the macroscopic level, the vast majority of volume is empty, with detailed representations only needed at object surfaces. Hence, using a full-size voxel grid to represent these structures is highly inefficient. To address this, we propose a hierarchical, coarse-to-fine approach for tokenizing 3D structures by leveraging their inherent sparsity. As shown in Fig. 1, our approach consists of two parts: (1) a hierarchical compression of 3D space using an octree, and (2) a fine-grained structural tokenization.
@@
The first component is the octree, an efficient data structure for losslessly compressing 3D grids. We begin with a single grid cell encompassing the entire 3D structure and recursively subdivide it based on a simple rule: if a cell is not empty (i.e., contains atoms), it is further partitioned. At each subdivision, each dimension is halved, yielding $82^{3}=$ 8 equally sized subcells (hence the name "octree"). This subdivision continues for a predefined number of levels, denoted by $L$. Let $c_{0}$ be the cell length at root level; then the cell length at the final level, $L-1$, is given by $c_{L-1}=c_{0}/2^{L-1}$.
@@
The second component is fine-grained tokenization of structural details. While the octree provides a coarse representation of non-empty regions, it lacks finer details within these regions, such as atom types and precise 3D coordinates in microscopic structures or surface details in macroscopic structures. While deeper octrees, as used in previous work [20], can capture finer details, they are computationally inefficient due to increased octree tokens. Instead, we adopt a more efficient approach inspired by the concept of 2D image patches in recent image transformer models [21]. In our method, we extend this idea to 3D structures by tokenizing the fine-grained structural details within the final-level non-empty regions, which we refer to as “3D patches”. A well-established technique in the image domain is vector quantized tokenization, as employed in VQ-VAE [22]. Similarly, we can quantize fine-grained 3D patches into discrete variables and perform autoregressive prediction on them. Alternatively, we can leverage recent patch-level diffusion loss strategies from MAR [23] to directly model continuous vector representations of 3D patches.
@@
Finally, we concatenate these tokens level by level. Beyond token content, we represent each token’s positional information using its tree level and the spatial coordinates of its cell center. For instance, the root cell is at level 0 with a center at $(c_{0}/2,c_{0}/2,c_{0}/2)$. During autoregressive prediction, since octree tokens are dynamically unfolded level by level, the positions of all tokens at the current level are known based on the predictions from previous level. This explicit knowledge of the token position is crucial, as autoregressive models predict only token content. In contrast, previous point-based tokenization lacks inherent positional information. As a result, previous models must infer both precise positions and contents simultaneously, making the task significantly more challenging.
2-Level Subtree Compression
@@
Although our octree tokenization avoids the cubic growth of cells, it remains inefficient for large 3D structures. At each level, there are at most $8N$ tokens (where $N$ is the number of non-empty cells at the final level), leading to a total of up to $8NL$ tokens over $L$ levels—approximately two orders of magnitude larger than $N$. To further reduce the token count, we introduce 2-level subtree compression, which merges a 2-level subtree into a single token. Each 2-level subtree consists of one parent node and 8 child nodes. Since the parent’s type is always 1, the entire subtree can be fully represented by the types of its 8 children, resulting in $2562^{8}=$ 256 possible states. This compression reduces the total token count by approximately a factor of 8, yielding at most $N(L-1)$ tokens. For positional information (i.e., cell center and level) of the compressed nodes, we retain the parent’s position.
Fine-Grained Atom Tokenization for Microscopic Structures
For the microscopic 3D structures focused on in this study, we adopt a simple fine-grained tokenization approach instead of the universal solution for various 3D structures. Specifically, by setting the final-level cell size $c_{L-1}$ sufficiently small (e.g., $0.24\text{\AA}$ in our experiments), we ensure that each cell contains at most one atom. Consequently, each fine-grained token corresponds to a single atom, represented by four variables: the atom type and its 3D coordinates within the cell. To further discretize in-cell positions, we use a predefined resolution $c_{r}$ (e.g., $0.01\text{\AA}$), converting continuous coordinates into integers ranging from 0 to $N_{p}=c_{L-1}/c_{r}$. Formally, we denote the content of the $i$-th token as $(t_{i},\bm{e}_{i})$, where $t_{i}$ is the token type and $\bm{e}_{i}=(e_{i}^{0},e_{i}^{1},e_{i}^{2})\in\{0,1,\dots,N_{p}-1\}^{3}$ represents its discretized in-cell 3D coordinates. Since octree tokens do not have in-cell positions, $\bm{e}_{i}$ is set to the default value $N_{p}/2$.
Figure 3: (a) The proposed masked next-token prediction mechanism effectively addresses the challenge of dynamically varying token positions. (b) Based on the duplicated masked token, Uni-3DAR unifies multi-frame 3D structural generation with token-level and structure-level understanding tasks in a single model, assigning each token a distinct role.
@@ 2.2 Masked Next-Token Prediction for Dynamic Token Positions
In traditional autoregressive models, token positions follow a fixed pattern. For instance, in text models, the token at the $i$-th position is always followed by the token at position $i+1$. Due to this static structure, the next token’s position can be directly inferred from the current token’s position, and explicit modeling of the next position is unnecessary.
In our proposed coarse-to-fine tokenization method for 3D structures, token positions are inherently dynamic and vary across samples. This variability makes it much more challenging to infer the next token’s position, so it is preferable to explicitly provide the next position to the model.
One straightforward approach, as suggested by previous work [20], is to encode both the current and the next token positions. However, we found that this method did not yield satisfactory results. In particular, the dynamic nature of the next token’s position introduces randomness that acts like noise, adversely affecting the current token’s representation. Moreover, the model must simultaneously learn to encode a token’s content based on its current position and predict the next token’s content based on a variable next position. This dual learning burden poses significant challenges and may lead to degraded performance.
Recently, many generative approaches based on masked prediction have emerged [24, 23]. At masked positions, these methods replace the token with a “[MASK]” symbol and predict its content based on the token’s current positional information. This approach sidesteps issues related to dynamic next positions and has proven effective in generation scenarios where the token order is randomized [23]. However, masked prediction typically relies on bi-directional attention, whereas our tokenization unfolds level-by-level in a uni-directional manner, making the direct application of bi-directional masked prediction challenging. Moreover, masked generation depends on parallel, independent sampling, which differs fundamentally from the causal sampling used in next-token prediction. Consequently, additional rule-based sampling strategies are often required during inference to balance performance and efficiency [23].
To address these challenges, we propose a simple solution that integrates masked prediction into the next-token prediction framework. Specifically, for each token, we duplicate it so that both copies share the same current position, and then replace the content of the first copy with a “[MASK]” token. Next-token prediction is then applied exclusively on the masked token. Notably, this process effectively performs masked prediction: the model receives a “[MASK]” input at that position and is tasked with predicting the token content. Compared with standard next-token prediction, our approach better handles the challenge of dynamic next positions; and compared with conventional masked prediction, it avoids the drawbacks of parallel independent sampling while retaining the benefits of causal sampling. Although this method doubles the sequence length, the performance gains justify the trade-off (see Sec. 3.8). Furthermore, as shown in Sec. 3.9, our efficiency optimizations enable masked next-token prediction with an inference speed reduction of only 15% to 30% compared to standard next-token prediction.

### v3

# Unified Cross-Scale 3D Generation and Understanding via Autoregressive Modeling

@@ Abstract
3D structure modeling is essential across scales, enabling applications from fluid simulation and 3D reconstruction to protein folding and molecular docking. Yet, despite shared 3D spatial patterns, current approaches remain fragmented, with models narrowly specialized for specific domains and unable to generalize across tasks or scales. We propose Uni-3DAR, a unified autoregressive framework for cross-scale 3D generation and understanding. At its core is a coarse-to-fine tokenizer based on octree data structures, which compresses diverse 3D structures into compact 1D token sequences. We further propose a two-level subtree compression strategy, which reduces the octree token sequence by up to 8x. To address the challenge of dynamically varying token positions introduced by compression, we introduce a masked next-token prediction strategy that ensures accurate positional modeling, significantly boosting model performance. Extensive experiments across multiple 3D generation and understanding tasks, including small molecules, proteins, polymers, crystals, and macroscopic 3D objects, validate its effectiveness and versatility. Notably, Uni-3DAR surpasses previous state-of-the-art diffusion models by a substantial margin, achieving up to 256% relative improvement while delivering inference speeds up to 21.8x faster.
@@ 1 Introduction
3D structure modeling underpins a wide range of real-world applications, spanning the planetary-scale dynamics of celestial bodies to the angstrom-scale arrangements of atoms and electrons. At the macroscopic level, it enables 3D object reconstruction, computational fluid dynamics simulations, and climate forecasting; at the microscopic level, it supports protein structure prediction (Jumper et al., 2021), crystal generation (Jiao et al., 2023), molecular dynamics (Wang et al., 2018a), and molecular docking (Alcaide et al., 2024).
Despite these shared spatial principles, 3D modeling tasks have largely evolved in silos. Models tailored for macroscopic structures fail to transfer to microscopic domains, and even applications at similar scales rarely generalize. For instance, a model designed for crystal generation cannot be directly applied to protein folding (Xie et al., 2021; Jiao et al., 2023). This fragmented development hinders data reuse and results in redundant, highly specialized models rather than a unified solution.
To overcome this fragmentation, we propose Uni-3DAR, a unified autoregressive framework for cross-scale 3D generation and understanding. At its core is a tokenizer that efficiently compresses diverse 3D structures into discrete 1D token sequences. Leveraging these compressed sequences, our autoregressive model unifies generative and understanding tasks within a single architecture.
The proposed tokenizer uses an octree data structure to compress the full-size 3D grid both losslessly and efficiently. As illustrated in Fig. 2 (a) and (b), we construct an octree by recursively subdividing the space up to a maximum depth of $L$. To adapt to data sparsity, branches corresponding to empty regions are pruned, resulting in a maximum of $8^{L-1}$ leaf grid cells (but most will be pruned due to sparsity). We then introduce a fine-grained tokenization that encodes details within each occupied leaf cell (we call it a “3D patch”), such as atomic types and precise coordinates for molecules, or more general VQVAE tokens (Van Den Oord et al., 2017). Concatenating these tokens level by level produces a hierarchical, coarse-to-fine 1D token sequence that effectively represents the 3D structure (fig.˜2(c)). Furthermore, we compress each two-level subtree (eight subcells) into a single 8-bit token instead of assigning an individual occupancy token to each node (Fig. 2 (d)). Since each subcell is binary (empty or not), grouping eight subcells yields $2^{8}=256$ distinct states, reducing the sequence length approximately 8x and converting 8 binary classifications into one 256-class task.
Figure 1: Uni-3DAR Overview. (a) A coarse-to-fine octree-based tokenizer converts 3D structures into 1D sequences (details in fig.˜2). The tokens are modeled by an autoregressive transformer trained with masked next-token prediction (details in fig.˜3) and can be optionally conditioned on cross-modal inputs (e.g., text, biological sequences, spectra). A single model supports single- and multi-frame generation as well as token- and structure-level understanding. (b) An example of octree from coarse level to fine level. Uni-3DAR generates tokens in a coarse-to-fine order: high-level occupancy tokens followed by level-0 tokens that capture local details (e.g., atom types and coordinates). The merits of octree over other 3D representations are discussed in Appendix˜A.
@@
However, the octree compression with empty tokens pruned disrupts the spatial mapping, meaning adjacent tokens no longer correspond to uniform intervals in the original 3D space. Unlike in 2D images with fixed patch positions, the model cannot reliably predict the next token without knowing its explicit target coordinates. We found that simply appending the next position to the current token yielded unsatisfactory results. To address this challenge, we propose a masked next-token prediction strategy. As illustrated in Fig. 3 (a), our method duplicates each token so that it appears twice with the same positional embedding. We then replace the first copy with a [MASK] token. The model still performs next-token prediction, but the prediction is made exclusively at the masked position. This setup ensures that the prediction is conditioned on the correct positional information of the intended target token, effectively resolving the issue of dynamic token positions. Although this approach doubles the sequence length, it achieves significant performance gains as validated in section˜D.2.
Uni-3DAR is built on several technical innovations: (1) a coarse-to-fine octree-based tokenization for efficient representation, (2) a 2-level subtree compression to reduce sequence length, (3) a unified fine-grained structural representation (for “3D patch”) to capture local details, and (4) a masked next-token prediction strategy to handle dynamic token positions, which enable our key contributions:
1. Unified Cross-Scale 3D Modeling. Leveraging the proposed coarse-to-fine tokenizer, Uni-3DAR can process a wide range of 3D structures, from macroscopic to microscopic scales.
2. Unified Generation and Understanding. Uni-3DAR seamlessly unifies 3D structural generation and understanding tasks within a single framework. As illustrated in Fig. 3 (b), different tasks use distinct tokens, ensuring clear separation without interference.
3. High Efficiency. Thanks to the octree and two-level subtree compression, Uni-3DAR represents 3D structures with far fewer tokens. For example, while O Pinheiro et al. (2023) requires $32^{3}=262,144$ tokens for a small molecule, Uni-3DAR needs only hundreds, and can scale to large proteins with thousands of atoms with deeper octree levels (section˜3.4). Moreover, section˜D.3 shows that Uni-3DAR is approximately 21.8x faster than prior diffusion-based models.
@@
4. High Accuracy. Extensive experiments across diverse tasks—including macroscopic 3D shape generation (table˜4), molecular (table˜1), crystal generation (table˜2), protein pocket prediction (table˜6), molecular docking (table˜7), and molecular pretraining (tables˜8 and 9)—demonstrate Uni-3DAR’s superior or competitive performance compared to existing methods. Notably, Uni-3DAR consistently outperforms diffusion-based models. Ablation studies (table˜12) highlight the benefits of unifying generation and understanding and validate the effectiveness of each component.
Figure 2: Overview of Uni-3DAR tokenization (illustrated in 2D using quadtree for clarity). (a) Adaptive coarse-to-fine subdivision of grid cells, where darker nodes indicate non-empty cells that can be further partitioned. (b) This partitioning process constructs an octree, providing a lossless compression of the full-size 3D grid. (c) Uni-3DAR’s tokenization consists of two components: hierarchical spatial compression via an octree and fine-grained structural tokenization. Each node’s position is determined by its tree level and cell center. (d) The proposed 2-level subtree compression reduces the octree tokens by 8x (4x in the illustrated quadtree).
@@ 2.1 Dynamic Coarse-to-Fine Tokenization for 3D Structures
3D structures are inherently sparse: at the microscopic scale, most space is empty except for scattered atoms; at the macroscopic level, detailed representations are only needed at object surfaces, with most volume remaining empty. Using a full-size voxel grid is thus highly inefficient. To address this, we propose a hierarchical, coarse-to-fine tokenization of 3D structures that exploits this sparsity. As shown in Fig. 1, our approach consists of two parts: (1) a hierarchical compression of 3D space using an octree, and (2) a fine-grained structural tokenization.
@@
The first component is the octree, an efficient data structure for lossless 3D grid compression. Starting with a single cell covering the entire structure, we recursively subdivide it: if a cell contains atoms, it is partitioned further. Each subdivision halves each dimension, producing $2^{3}=8$ equal subcells (hence “octree”). This process continues for $L$ levels. If $c_{0}$ is the root cell length, the cell length at level L1L-1 is $c_{L-1}=c_{0}/2^{L-1}$. We refer to these leaf subcells as fine-grained "3D patches," which are then tokenized as detailed in the following paragraph.
@@
The second component is the fine-grained tokenization of structural details. While the octree effectively identifies coarse, non-empty regions, it lacks finer details such as atom types and precise coordinates (microscopic) or surface features (macroscopic). Although using deeper octrees can capture more detail (Ibing et al., 2023), this approach becomes inefficient due to the rapidly increasing number of tokens. Instead, inspired by the use of 2D image patches (Alexey, 2020), we treat the contents of each final-level non-empty region as a “3D patch.” These patches can be processed in various ways; for instance, they can be quantized into discrete tokens for autoregressive prediction, similar to VQ-VAE (Van Den Oord et al., 2017), or modeled using a patch-level diffusion loss for continuous vector representations (Li et al., 2024b) (ablation in Table˜12). In our experiments, we demonstrate this flexibility by using raw atom types and coordinates as fine-grained tokens for microscopic data (we set the patch size to ensure each 3D patch only contains one atom), and VQ-VAE tokens for macroscopic data. More details are in Appendix˜B.
@@
Finally, we concatenate tokens level by level. Beyond token content, we represent each token’s positional information using its tree level and the spatial coordinates of its cell center. For instance, the root cell is at level 0 with a center at $(c_{0}/2,c_{0}/2,c_{0}/2)$. During autoregressive prediction, since octree tokens are dynamically unfolded level by level, the positions of all tokens at the current level are known based on the predictions from previous level. This explicit knowledge of the token position is crucial, as autoregressive models predict only token content.
@@
2-Level Subtree Compression    Although octree tokenization avoids cubic cell growth, it remains inefficient for large 3D structures. Each level has up to $8N$ tokens ($N$ = non-empty final-level cells), totaling up to $8NL$ tokens across $L$ levels, about two orders of magnitude larger than $N$. To reduce this, we introduce 2-level subtree compression, merging a parent and its 8 children into a single token. As the parent’s type is always 1, the subtree is fully represented by its 8 children’s types, yielding $2^{8}=256$ possible states. This cuts token count by  8×, down to at most N(L1)N(L-1) tokens. For positional information of the compressed nodes, we retain the their parent’s center and level.
Figure 3: (a) Masked Next-Token Prediction. To handle the challenge of dynamically positioned tokens in sparse 3D structures, Uni-3DAR decouples position and content generation. Unlike standard next-token prediction, we first infer the next token’s position from the octree hierarchy, place a “[MASK]” token, and then have the model predict only its content (e.g., occupancy or fine-grained properties). (b) Unified Framework for 3D Generation and Understanding. The Uni-3DAR architecture is a versatile, multi-task model. It supports autoregressive generation of complex 3D structures (blue arrows) and can be prompted to perform both token-level (green arrows) and structure-level (blue box) understanding tasks within a single framework.
@@ 2.2 Masked Next-Token Prediction for Dynamic Token Positions
In standard autoregressive models, such as those for text, token positions follow a fixed, sequential order (e.g., token $i+1$ always follows token $i$). This static structure makes the next token’s position implicitly known, obviating the need for its explicit prediction. In contrast, our coarse-to-fine 3D tokenization generates a token sequence where positions are dynamic and vary across different structures. This variability introduces a significant challenge: inferring the next token’s position becomes non-trivial, making it preferable to provide this information to the model explicitly.
A straightforward approach is to encode both the current and next positions within each token (Ibing et al., 2023). However, we found this method leads to suboptimal performance (Table˜12). We hypothesize that the unpredictable nature of the next token’s position introduces noise that degrades the current token’s representation. This intuition is supported by prior work like Yan et al. (2022), which decouples position and content prediction into two separate transformer modules.
Another promising direction is masked prediction (Chang et al., 2022; Li et al., 2024b), where a model predicts the content of a masked token given its position. This has proven effective for generative tasks with non-sequential or randomized token orders (Li et al., 2024b; Pang et al., 2024). However, directly applying conventional masked prediction to our framework is problematic. First, it typically relies on bi-directional attention, whereas our hierarchical tokenization unfolds uni-directionally. Second, it often requires parallel, non-causal sampling, which necessitates complex, rule-based inference strategies to balance performance and efficiency (Li et al., 2024b).
To resolve these issues, we introduce Masked Next-Token Prediction (MNTP), a simple yet effective method that integrates masked prediction into a standard autoregressive framework. The core idea is to duplicate each token. For a given token, we first generate a placeholder with its content replaced by a special [MASK] symbol while retaining its correct position. This is immediately followed by a second token at the same position but with the true content. The model’s objective is to predict the content of this second token, conditioned on the [MASK] token and all preceding tokens.
This formulation effectively reframes next-token prediction as a masked prediction task: the model is given a position with a mask and asked to fill in the content. This approach elegantly sidesteps the challenge of predicting dynamic next positions. Compared to conventional masked prediction, MNTP preserves a strictly causal, uni-directional attention flow, eliminating the need for complex sampling schemes. While this duplication doubles the sequence length, we demonstrate in Section˜D.2 that the substantial performance gains justify this trade-off. Furthermore, through targeted optimizations discussed in Section˜D.3, the impact on inference latency is modest, with only a 15%–30% increase compared to standard next-token prediction (fig.˜4).